# SPACE: SPike-Aware Consistency Enhancement for Test-Time Adaptation in Spiking Neural Networks

### v2

@@ Appendix A Kernel-Based Consistency Regularization
To investigate whether higher-dimensional feature relationships can enhance the consistency measure, we integrate kernel embeddings into the loss function. The channel-wise feature probability distribution $\mathbf{P}_{c}(\mathbf{x}_{i})$ can be mapped into a higher-dimensional reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, using a Gaussian kernel defined as
@@
where $\sigma$ is the kernel bandwidth. The kernel function $k(\mathbf{z},\mathbf{z}^{\prime})$ implicitly defines a mapping $\phi$: $\mathcal{Z}\xrightarrow{}\mathcal{H}$, such that the inner product in $\mathcal{H}$ is given by the kernel:
@@
$\begin{split}\text{MMD}^{2}(\mathbf{P}_{c}(\mathbf{x}_{i}),\mathbf{P}_{c}(\mathbf{x}_{j}))=\frac{1}{D^{2}}\|\sum^{D}_{d=1}\phi(\mathbf{z}_{i_{d}})-\sum^{D}_{d=1}\phi(\mathbf{z}_{j_{d}})\|^{2}_{\mathcal{H}}.\end{split}$ (11)
Then we average the MMD values across all channels:
@@
$\mathcal{L}^{*}(\theta;\mathbf{x})\triangleq\mathcal{L}+\lambda_{\text{MMD}}\sum_{1\leq j<i\leq M}\text{MMD}^{2}(i,j|\mathbf{x})$ (13)
To evaluate the impact of incorporating kernel embedding into the loss function, we tested its performance on CIFAR-10-C dataset, using SNN-VGG9 model, as shown in Table 3. The results indicate that adding kernel embedding provides little to no improvement over the original SPACE method, with the average performance remaining nearly identical. This limited effectiveness can be attributed to several factors. First, the distribution differences introduced by augmentations, such as noise, blur, and weather effects, are relatively small, reducing the need for advanced alignment mechanisms like kernel embedding. Additionally, the existing similarity measures in SPACE already effectively capture the relationships between augmented samples, leaving little room for further optimization. Moreover, the added computational complexity of kernel embedding, especially the cost of kernel calculations, may even slightly hinder performance in scenarios where the original approach is already sufficient. These factors combined suggest that kernel embedding is not particularly beneficial in this experimental setting.
@@
Appendix B Exploration of Alternative Alignment Objectives
B.1 Two Alternative Alignment Objectives
Average Membrane Potential (SPACE + a.m.p.)
In the main paper, we use feature maps based on the spike counts over the entire time window as the alignment objective. In this section, we explore an alternative biologically inspired metric: the average membrane potential obtained from the LIF neurons in the final layer of the feature extractor. Specifically, for a given test point $\mathbf{x}_{i}$, the membrane potentials are captured in the feature map $\mathbf{U}(\mathbf{x}_{i})\in\mathbb{R}^{T\times C\times D}$, where $T$, $C$, and $D$ represent the time steps, the number of channels, and the spatial dimensions, respectively. For LIF neurons that spike at a specific time step, we use the membrane potential after the LIF reset, which indirectly captures the spiking information. To obtain the primary objective feature map $\mathbf{F}(\mathbf{x}_{i})$, we compute the temporal average of $\mathbf{U}(\mathbf{x}_{i})$ along the first dimension. The resulting feature map is then used to calculate the loss following Equations 4 and Equation 5 in the main paper. Finally, the model is updated via SGD.
Spikes through All Time Steps (SPACE + s.t.t.)
Instead of summing the spike counts over the temporal dimension, we explicitly retain the temporal information, resulting in a feature map with an additional dimension that captures the spiking activity at each time step. This approach preserves the precise temporal dynamics of the spikes and leverages them directly for alignment. Specifically, we align the channel-wise spiking activity across all time steps between each pair of augmented samples. The update rules remain consistent with those described in the main paper.
B.2 Experimental Results and Analysis
@@
We implement the two approaches above and evaluate on CIFAR-10-C dataset with SNN-VGG9 model. The results are shown in Table 4.
Compared to No Adapt, SPACE + a.m.p. improves the generalization ability of the model. However, its performance is noticeably inferior to SPACE on the majority of corruptions. This may be due to the fact that the membrane potentials of the same sample under different augmentations tend to exhibit similar patterns, making them less effective in capturing the temporal distribution characteristics of SNNs. For example, in extreme cases where the membrane potential of a LIF neuron stays near the firing threshold throughout all time steps without crossing it, SPACE + a.m.p. fails to capture this information. In contrast, SPACE can reflect this behavior through spike counts, which more directly encode spiking activity.
SPACE + s.t.t. fails to improve the generalization ability of the model. We hypothesize that this is due to the model overfitting during the alignment process. Specifically, the updates made to minimize the alignment loss may inadvertently lead to excessive specialization to the augmented samples, thereby hindering its ability to generalize effectively to unseen corruptions.
Appendix C Additional Ablation Study
Effect of Augmentation Quantity per Test Point
In the main paper, we follow MEMO [47] to determine the number of augmentations $M$. To analyze the effect of $M$ on our method, we evaluated the average accuracy and the evaluation time per test point the same computing resource conditions on CIFAR-10-C using SNN-VGG9 with varying augmentation quantities $M=\{2,4,8,16,32,64,128\}$. As shown in Table 5, at least 8 augmentations are required to achieve noticeable performance improvement. For $M=\{32,64,128\}$, while the performance continues to improve gradually, the evaluation time increases significantly. Considering the trade-off between accuracy and evaluation time, $M=32$ is the optimal choice.
No Adapt Number of Augmentation
2 4 8 16 32 64 128
Accuracy 66.57 54.60 66.91 69.77 70.60 71.03 71.07 71.11
Evaluation Time 0.144 0.318 0.320 0.327 0.339 0.345 0.443 0.788
Appendix D Additional Experimental Information
D.1 More Details on Datasets
In this paper, we conducted six static datasets and one neuromorphic dataset to evaluate the out-of-distribution generalization ability. They are CIFAR-10-C, CIFAR-100-C, Tiny-ImageNet-C [15], ImageNet-V2 [35], ImageNet-R [17], ImageNet-A [18] and DVS Gesture-C [20].
CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C
ImageNet V2/R/A
ImageNet-V2, ImageNet-R, and ImageNet-A are three datasets designed to evaluate model robustness under different real-world challenges. ImageNet-V2 consists of a new set of images collected under similar conditions as the original ImageNet [6], serving as a test of distribution shift. ImageNet-R focuses on image classification robustness across various artistic renditions, such as paintings, sketches, and cartoons, covering 200 ImageNet classes. ImageNet-A, on the other hand, is an adversarially filtered subset of natural images that models often misclassify, targeting their inherent weaknesses. Compared to corruption-based datasets, these benchmarks provide a better reflection of real-world scenarios and challenges.
DVS Gesture-C
DVS Gesture-C is a corrupted variant of the standard DVS Gesture [1] dataset. This variant introduces six distinct corruption types: DropPixel, DropEvent, RefractoryPeriod, TimeJitter, SpatialJitter, and UniformNoise. These corruptions, implemented via the Tonic API [29], effectively simulate real-world imperfections in event-based data, including sensor noise and timing inaccuracies.
D.2 More Details on Pretrained Models
SNN-VGG with BNTT
We adopt SNN-VGG with Batch-Normalization Through Time (BNTT) [24] as our main backbone model. BNTT decouples the parameters along the time axis within each layer, effectively capturing the temporal dynamics of spikes. The temporally evolving learnable parameters in BNTT enable neurons to regulate their spike rates across different time steps, facilitating low-latency and low-energy training from scratch. The pre-trained model achieves accuracies of 90.61%, 69.50% and 58.36% on the original CIFAR-10, CIFAR-100 and Tiny-ImageNet test sets respectively.
SNN-ResNet
Spike-driven Transformer V3
SNN-ConvLSTM
DVS Gesture is a neuromorphic dataset with a temporal dimension, making it well-suited for evaluating RNN-based architectures. To validate the effectiveness of our work on such networks, we employed a spike-driven 2ConvLSTM-1FC model, with hidden layer dimensions of 32 and 64, respectively. This network achieves an accuracy of 93.36% on the original DVS Gesture test set.
D.3 More Details on Implementation
To better reflect real-world scenarios, we adapt our method to independently perform adaptation on each individual test input. In the same experiment, the learning rate remains fixed across different shifted domains, as the domain of a given test sample is typically unknown in real-world settings. For each set of experiments, we employ the SGD optimizer and evaluate learning rates ranging from 0.001 to 0.8, selecting the optimal value based on performance. We trained and tested all models and datasets in 4 NVIDIA GeForce RTX 3090 GPUs.

### v3

@@ Appendix B Kernel-Based Consistency Regularization
To investigate whether higher-dimensional feature relationships can enhance the consistency measure, we integrate kernel embeddings into the loss function. The channel-wise feature probability distribution $\mathbf{P}$ can be mapped into a higher-dimensional reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, using a Gaussian kernel defined as
@@
where $\sigma$ is the kernel bandwidth. The kernel function $k(\mathbf{z},\mathbf{z}^{\prime})$ implicitly defines a mapping $\phi$: $\mathcal{Z}\xrightarrow{}\mathcal{H}$, such that the inner product in $\mathcal{H}$ can be given by the kernel function:
@@
$\text{MMD}^{2}(\mathbf{P}_{c}(\mathbf{x}_{i}),\mathbf{P}_{c}(\mathbf{x}_{j}))=\frac{1}{D^{2}}\Bigg\|\sum^{D}_{d=1}\phi(\mathbf{z}_{i_{d}})-\sum^{D}_{d=1}\phi(\mathbf{z}_{j_{d}})\Bigg\|^{2}_{\mathcal{H}}.$ (11)
Then we average the MMD values across channels:
@@
$\mathcal{L}^{*}(E_{\theta_{E}};\mathbf{x})\triangleq\mathcal{L}+\lambda_{\text{MMD}}\sum_{1\leq j<i\leq M}\text{MMD}^{2}(i,j|\mathbf{x})$ (13)
We evaluated augmenting SPACE with a kernel embedding loss (MMD) on CIFAR-10-C using SNN-VGG9 (Table 6). The average accuracy is virtually unchanged and occasionally slightly worse than our base alignment. We hypothesize three reasons. 1) Representation sufficiency: the spike-count feature map, coupled with our linear, channel-wise similarity, already captures the essential distributional alignment across augmentations; mapping to an RKHS adds little signal. 2) Regime mismatch: our augmentations are semantics-preserving and bounded in severity (e.g., AugMix [19]), inducing small, approximately uni-modal shifts where MMD—often helpful under large or multi-modal discrepancies—provides limited additional benefit. 3) Estimation and cost: per-sample TTA affords only a small augmented batch, making MMD estimates (and their gradients) higher-variance and bandwidth-sensitive, while adding nontrivial compute. These factors combined suggest that kernel embedding is not particularly beneficial in this experimental setting.
Appendix C Additional Experimental Information
C.1 Details on Datasets
In this paper, we conducted six static datasets and one neuromorphic dataset to evaluate the out-of-distribution generalization ability. They are CIFAR-10-C [^1] https://zenodo.org/records/2535967, CIFAR-100-C [^2] https://zenodo.org/records/3555552, Tiny-ImageNet-C [^3] https://zenodo.org/records/2469796 [18], ImageNet-V2444https://huggingface.co/datasets/vaishaal/ImageNetV2/tree/main [42], ImageNet-R [^5] https://github.com/hendrycks/imagenet-r [20], ImageNet-A [^6] https://github.com/hendrycks/natural-adv-examples [21] and DVS Gesture-C [^7] http://research.ibm.com/dvsgesture/ [23].
CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C
ImageNet-V2, ImageNet-R and ImageNet-A
These three datasets are designed to evaluate model robustness under different real-world challenges. ImageNet-V2 consists of a new set of images collected under similar conditions as the original ImageNet [7], serving as a test of distribution shift. ImageNet-R focuses on image classification robustness across various artistic renditions, such as paintings, sketches, and cartoons, covering 200 ImageNet classes. ImageNet-A, on the other hand, is an adversarially filtered subset of natural images that models often misclassify, targeting their inherent weaknesses. Compared to corruption-based datasets, these benchmarks provide a better reflection of real-world scenarios and challenges.
DVS Gesture-C
DVS Gesture-C is a corrupted variant of the standard DVS Gesture [1] dataset. This variant introduces six corruption types: DropPixel, DropEvent, RefractoryPeriod, TimeJitter, SpatialJitter, and UniformNoise. These corruptions, implemented via the Tonic API [35], effectively simulate real-world imperfections in event-based data, including sensor noise and timing inaccuracies.
C.2 Details on Pretrained Models
SNN-VGG with BNTT
 [^8] https://github.com/Intelligent-Computing-Lab-Yale/BNTT-Batch-Normalization-Through-Time
We adopt SNN-VGG with Batch-Normalization Through Time (BNTT) [29] as our main backbone model. BNTT decouples parameters along the time axis within each layer, effectively capturing spikes’ temporal dynamics. The temporally evolving parameters in BNTT enable neurons to regulate spike rates across time steps, facilitating low-latency, low-energy training from scratch. The pre-trained model achieves accuracies of 90.61%, 69.50%, and 58.36% on the CIFAR-10, CIFAR-100, and Tiny-ImageNet test sets, respectively.
SNN-ResNet
 [^9] https://github.com/chan8972/Enabling_Spikebased_Backpropagation
Spike-driven Transformer V3
 [^10] https://github.com/BICLab/Spike-Driven-Transformer-V3
SNN-ConvLSTM
DVS Gesture is a neuromorphic dataset with a temporal dimension, making it well-suited for evaluating RNN-based architectures. To validate the effectiveness of our work on such networks, we employed a spike-driven 2ConvLSTM-1FC model, with hidden layer dimensions of 32 and 64. This network achieves an accuracy of 93.36% on the original DVS Gesture test set.
C.3 Details on Implementation
To better reflect real-world scenarios, we adapt our method to independently perform adaptation on each individual test input. In the same experiment, the learning rate remains fixed across different shifted domains, as the domain of a given test sample is typically unknown in real-world settings. For each set of experiments, we employ the SGD optimizer and evaluate $\eta$ ranging from 0.001 to 0.8, selecting the optimal value based on performance. We trained and tested all models and datasets in 4 NVIDIA GeForce RTX 3090 GPUs.
Appendix D Additional Ablation Studies
Table 7: Performance comparison of SPACE with different augmentation strength on CIFAR-10-C (Gaussian Noise) with SNN-VGG9 model regarding Accuracy (%).
Strength $s$ 1 3 5 7 10 Larger Boundary
Accuracy 77.98 77.88 77.48 77.41 77.02 74.52
Effect of Augmentation Strength
We study the impact of AugMix [19] strength $s$ on SPACE. Following MEMO, our default is s=1s{=}1; we sweep $s\in[1,10]$ while keeping all other AugMix settings fixed (mixture width/depth, operators) and respecting AugMix’s semantic bounds (e.g., rotation $\leq 30^{\circ}$, translation $\leq\pm 16$ px). As shown in Table 7, accuracy remains stable within these bounds, with only minor declines as $s$ increases. When we deliberately exceed the bounds (e.g., rotation $50^{\circ}$, translation $\pm 18$ px), accuracy drops noticeably, indicating that overly aggressive transformations inject harmful noise and destabilize per-sample alignment under the augmented batches. In this regime, the alignment objective tends to match augmentation-induced artifacts (noise) rather than invariant, class-relevant structure in the feature maps, thereby undermining TTA and diminishing its gains. These results are consistent with AugMix’s design goal of bounded, semantics-preserving perturbations; we therefore adopt s=1s{=}1 by default and recommend staying within AugMix’s prescribed ranges for robust adaptation.
Effect of Encoding Methods for SNNs
We assess sensitivity to spike encoding by replacing Poisson rate coding with temporal coding in SNN-VGG9 while keeping architecture, training, and TTA hyperparameters fixed. As reported in Figure 5, the temporal-coded baseline is markedly less robust under domain shift than its rate-coded counterpart. This aligns with intuition: temporal coding relies on precise spike times and is therefore vulnerable to perturbation-induced timing jitter, whereas rate coding averages over spikes, providing redundancy and noise tolerance.
Figure 5: Performance of SNNs with different encoding under domain shift. Values are Top-1 accuracy (%); Acc. Loss = Clean - Shifted.
Encoding Method Clean Shifted Acc. Loss
Temporal Coding 89.27 58.11 31.16
Rate Coding 90.61 66.57 24.04
Table 8 shows that SPACE improves both models, with consistently larger relative gains for the temporal-coded network. The mechanism is a temporal-boundary effect: the total spike count is a sensitive proxy for timing stability because small perturbations can shift spikes across the fixed processing window; for spikes near the boundary, tiny jitters flip their contribution from 1 to 0, causing discrete changes in the count. By enforcing consistency of total counts across augmentations, SPACE penalizes these boundary-crossing events and implicitly forces the model to stabilizes spike timing without brittle spike-to-spike matching. The resulting gradients are stronger in temporal-coded models than in rate-coded ones, yielding larger robustness gains.
Table 8: Performance comparison of spike encoding methods for SNN-VGG9 model on CIFAR-10-C regarding Accuracy (%). The bold number indicates the performance improvement.
@@ Appendix D Additional Ablation Studies
Temporal Coding
No Adapt 39.74 39.32 68.42 59.59 54.43 60.19 65.90 69.78 46.21 66.77 78.14 27.26 70.29 46.00 79.59 58.11
SPACE 45.19 44.00 71.82 64.16 56.31 64.90 68.65 70.29 56.90 72.40 79.09 68.64 70.66 50.06 80.07 64.21 (+6.10)
Rate Coding
@@
SPACE 77.98 79.34 69.41 71.59 67.76 72.14 74.67 78.43 52.80 79.59 83.22 23.85 75.49 76.24 82.88 71.03 (+4.46)
@@
Exploration of Alternative Alignment Objectives
We evaluate two alternatives to SPACE’s default count-based feature map on CIFAR-10-C with SNN-VGG9, keeping architecture, training, and TTA hyperparameters fixed. Results are summarized in Table 9.
1) Average Membrane Potential (SPACE + a.m.p.). Let $\mathbf{U}(\mathbf{x})\in\mathbb{R}^{T\times C\times D}$ denote the LIF membrane traces from the selected layer (see subsection 3.2). For spikes, the post-reset potential is used. We form $\mathbf{F}(\mathbf{x})=\tfrac{1}{T}\sum_{t=1}^{T}\mathbf{U}_{t}(\mathbf{x})\in\mathbb{R}^{C\times D}$ and apply the same channel-wise consistency loss as SPACE. It improves over No Adapt but is consistently worse than count-based SPACE on most corruptions. Membrane traces are smooth and highly correlated across augmentations, making them less sensitive to discrete spiking changes; near-threshold, non-firing trajectories can be indistinguishable in $\mathbf{F}(\mathbf{x})$, whereas counts capture such events directly.
2) Spikes Through Time (SPACE + s.t.t.). Instead of summing over time, we retain the temporal dimension of the spike tensor $\mathbf{S}(\mathbf{x})\in\mathbb{R}^{T\times C\times D}$ and align per-channel activity at each time step across augmentations (loss averaged over $t$); optimization is unchanged. This variant does not improve robustness. Keeping the full temporal waveform inflates variance and forces the model to match augmentation-specific transients: benign timing jitter (e.g., a spike at t=10t{=}10 vs. t=12t{=}12) is penalized as a mismatch, producing conflicting gradients that push spike times to align across views rather than stabilizing the underlying excitability. The model thus overfits to view-specific temporal shapes, hurting generalization, while also incurring higher compute and memory. We further applied Gaussian smoothing along time before alignment, $\tilde{\mathbf{S}}=g_{\sigma}*\mathbf{S}$, to reduce jitter-induced overfitting. Results remained suboptimal. With small $\sigma$, smoothing is too weak to suppress shift-induced conflicts; with larger $\sigma$, it erases informative peaks and attenuates counts. The core issue is not merely information loss but the learning signal each representation provides: spike counts yield a shift-invariant, macro-level target that is robust to timing jitter and directs clear adjustments to neuron excitability, whereas smoothed sequences still demand matching temporal shapes, preserving shift-sensitivity and inducing noisy, contradictory gradients.
Consequently, SPACE’s count-based objective offers a more stable and effective alignment signal for our TTA setting.
Table 10: Wilcoxon signed-rank test on CIFAR-10-C, CIFAR-100-C, and Tiny-ImageNet-C with CNN-based SNN models. W denotes the Wilcoxon signed-rank statistic and p is the two-sided p-value.
Compared Method Metric Noise Blur Weather Digital Average
CIFAR-10-C, SNN-VGG9
vs. No Adapt W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 0 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 1.0000 0.0312
vs. RoTTA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 0 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 1.0000 0.0312
vs. DeYO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. SITA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 0 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 1.0000 0.0312
vs. MEMO W 15 0 15 15 15 15 15 15 15 15 10 9 0 15 0 15
p 1.0000 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.4000 0.3333 0.0312 0.0312 1.0000 0.0312
CIFAR-100-C, SNN-VGG9
vs. No Adapt W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. RoTTA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. DeYO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. SITA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. MEMO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 10 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.4000 0.0312
Tiny-ImageNet-C, SNN-VGG11
vs. No Adapt W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. RoTTA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. DeYO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. SITA W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. MEMO W 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
Table 11: Wilcoxon signed-rank test at the per-class level on CIFAR-10-C (Gaussian/Shot Noise) with CNN-based SNN models. W denotes the Wilcoxon signed-rank statistic and p is the two-sided p-value.
Compared Method Metric Class Index
0 1 2 3 4 5 6 7 8 9
Gaussian Noise
vs. No Adapt W 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. RoTTA W 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. DeYO W 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. SITA W 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. MEMO W 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
Shot Noise
vs. No Adapt W 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. RoTTA W 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. DeYO W 15 15 15 15 15 15 15 15 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312
vs. SITA W 15 15 15 15 15 15 15 14 15 15
p 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0312 0.0625 0.0312 0.0312
vs. MEMO W 5 0 0 5 5 0 0 0 14 0
p 0.7812 1.0000 1.0000 0.7812 0.7812 1.0000 1.0000 1.0000 0.0625 1.0000
Statistical Significance
We assess statistical significance using the Wilcoxon signed-rank test. Specifically, for each of the three main experiments, we run five independent trials and apply a paired Wilcoxon test between SPACE and each baseline across the five runs; reported $p$-values in Table 10 confirm that SPACE’s gains are statistically significant. To further strengthen the evidence, we select Gaussian and Shot Noise from CIFAR-10-C and conduct Wilcoxon signed-rank tests at the per-class level (Table 11), demonstrating that the improvements are consistent across categories rather than driven by a few classes.